Latin Corporation’s breakeven point in revenues is $1,000,000 and fixed costs is $400,000Required: 1)Compute the contribution margin ratio2)Compute the selling price if variable costs are $12 per unit3)Suppose 80,000 units are sold, compute the margin of safety


1
Expert's answer
2022-01-16T17:50:55-0500

1) The contribution margin ratio is:

400,000/1,000,000×100% = 40%.

2) The selling price if variable costs are $12 per unit is:

Q = VC/AVC = 600,000/12 = 50,000 units,

400,000/(P - 12) = 50,000,

P - 12 = 8,

P = 20.

3) If 80,000 units are sold, then the margin of safety is:

(80,000 - 50,000)/80,000×100% = 37.5%.